Determine (input the letter only) what functions does the following expression belong.

The questions items has the format of f(x), f(m, n).

A: one to one, not onto B: onto, not one to one

C: one to one, onto D: Neither one to one nor onto

E: Not a function

Given for solving the f(S): f(x) = x^2/3 

The questions items has the format of S={1,2,3}.

Format of answer: f(S) = {1, 2, 3} 


  1. f(x) = x + 2, Z → Z.
  2. f(x) = 4x+5, R → R.
  3. For the given sequence, provide a simple formula or rule that generates the terms of the sequence: 2, 3, 7, 25, 121, 721, 5041, 40321,..
  4. f(x) = 2x + 1, R→ R.
  5. f(x) = x^3 , R → R.
  6. f(x)= x + 2 / x^3 + 8, Z → Z,
  7. S = {1, 5, 7, 11}
  8. S = {2, 6, 10, 14}.
1
Expert's answer
2021-06-01T03:45:59-0400
  1. f is neither one-to-one nor onto
  2. f(x) = x + 2, Z → Z.
  3. 3,5,8,12,17,23,30,38,47,... The sequence starts with 3, then 2 is added to the previous term, then 3 is added to the previous term, then 4 is added to the previous term, and so on. We note that 47 is 9 added to 38, thus the next three terms add 10, then 11 and then 12. aio=ag + 10 = 47 + 10 = 57 all =a10+11.57+11.68 ail=all+12=68+12=80